# Policy Mirror Descent with Lookahead

### v1

@@ 1 Introduction
Policy Mirror Descent (PMD) is a general class of algorithms for solving reinforcement learning (RL) problems. Motivated by the surge of interest in understanding popular Policy Gradient (PG) methods, PMD has been recently investigated in a line of works (Lan, 2023; Xiao, 2022; Li et al., 2023; Johnson et al., 2023). Notably, PMD encompasses several PG methods as particular cases via its flexible mirror map. A prominent example is the celebrated Natural PG method. PMD has also close connections to state-of-the-art methods such as TRPO and PPO (Tomar et al., 2022) which have achieved widespread empirical success (Schulman et al., 2017a, b), including most recently in fine-tuning Large Language Models via RL from human feedback (Ouyang et al., 2022). Interestingly, PMD has also been inspired by one of the most fundamental algorithms in RL: Policy Iteration (PI). While PI alternates between policy evaluation and policy improvement, PMD regularizes the latter step to address the instability issue of PI with inexact policy evaluation.
@@
  • We examine the inexact setting where the $h$-step action value functions are not available. In this setting, we propose a Monte Carlo sample-based procedure to estimate the unknown $h$-step lookahead action-value function at each state-action pair.
    We provide a sample complexity for this Monte Carlo procedure, improving over standard PMD thanks to the use of lookahead. Larger lookahead depth translates into a better sample complexity and a faster suboptimality gap convergence rate. However, this improvement comes at a more intensive computational effort to perform the lookahead steps. Sec. 5 discusses this inexact setting and the aforementioned tradeoff.
  • We extend our results to the function approximation setting to address the case of large state spaces where tabular methods are not tractable. In particular, we design a $h$-PMD algorithm where the $h$-step lookahead action-value function is approximated using a linear combination of state-action feature vectors in the policy evaluation step. Under linear function approximation and using a generative sampling model, we provide a performance bound for our $h$-PMD algorithm with linear function approximation. Our resulting sampling complexity only depends on the dimension of the feature map space instead of the size of the state space and improves over prior work analyzing PMD with function approximation.
@@ 2 Preliminaries
Notation. For any integer $n\in\mathbb{N}\setminus\{0\}$, the set of integers from $1$ to $n$ is denoted by $[n]$. For a finite set $\mathcal{X},$ we denote by $\Delta(\mathcal{X})$ the probability simplex over the set $\mathcal{X}\,.$ For any integer $d\in\mathbb{N}$, we endow the Euclidean space ${{\mathbb{R}}}^{d}$ with the standard maximum norm $\|\cdot\|_{\infty}$ defined for every $v\in{{\mathbb{R}}}^{d}$ by $\|v\|_{\infty}:=\max_{i\in[d]}|v_{i}|$ where $v_{i}$ is the $i$-th coordinate. The relative interior of a set $\mathcal{Z}$ is denoted by $\operatorname{ri}(\mathcal{Z})\,.$
@@
The Bellman optimality operator ${\mathcal{T}}$ and the expected Bellman operator ${\mathcal{T}^{\pi}}$ for any given policy $\pi\in\Pi$ are $\gamma$-contraction mappings w.r.t. the $\|\cdot\|_{\infty}$ norm and hence admit unique fixed points $V^{\star}$ and $V^{\pi}$ respectively. In particular, these operators satisfy the following inequalities: $\lVert{\mathcal{T}^{\pi}}V-V^{\pi}\rVert_{\infty}\leq\gamma\lVert V-V^{\pi}\rVert_{\infty}$ and $\lVert{\mathcal{T}}V-V^{\star}\rVert_{\infty}\leq\gamma\lVert V-V^{\star}\rVert_{\infty}$ for any $V\in{{\mathbb{R}}}^{S}\,.$ Moreover, we recall that for $\mathbf{e}=(1,1,\ldots,1)\in\mathbb{R}^{S}$ and any $c\in\mathbb{R}_{\geq 0},V\in{{\mathbb{R}}}^{S}$, we have ${\mathcal{T}^{\pi}}(V+ce)={\mathcal{T}^{\pi}}V+\gamma c\,\mathbf{e}$ and ${\mathcal{T}}(V+ce)={\mathcal{T}}V+\gamma c\,\mathbf{e}\,.$ We recall that the set $\mathcal{G}(V^{\star}):=\{\pi\in\Pi:{\mathcal{T}^{\pi}}V^{\star}={\mathcal{T}}V^{\star}\}$ of $1$step greedy policies w.r.t. the optimal value function $V^{\star}$ coincides with the set of stationary optimal policies.
@@ 3.1 PMD and its connection to PI
Policy Gradient methods consist in performing gradient ascent updates with respect to an objective $\mathbb{E}_{\rho}[V^{\pi}(s)]$ w.r.t. the policy $\pi$ parametrized by its values $\pi(a|s)$ for $s,a\in\mathcal{S}\times\mathcal{A}$ in the case of a tabular policy parametrization. Interestingly, policy gradient methods have been recently shown to be closely connected to a class of Policy Mirror Descent algorithms by using dynamical reweighting of the Bregman divergence with discounted visitation distribution coefficients in a classical mirror descent update rule using policy gradients. We refer the reader to section 4 in Xiao (2022) or section 3.1 in Johnson et al. (2023) for a detailed exposition.
The resulting PMD update rule is given by
@@
where we use the shorthand notations $\pi^{k}_{s}={\pi_{k}}(\cdot|s)$ and $\pi_{s}=\pi(\cdot|s)$, where $(\eta_{k})$ is a sequence of positive step sizes, $Q^{\pi_{k}}_{s}\in\mathbb{R}^{A}$ is a vector containing state action values for ${\pi_{k}}$ at state $s$, and $D_{\phi}$ is the Bregman divergence induced by a mirror map $\phi:\operatorname{dom}\phi\to{{\mathbb{R}}}$ such that $\Pi\subset\operatorname{dom}\phi\,,$ i.e. for any $\pi,\pi^{\prime}\in\Pi,$
$D_{\phi}(\pi,\pi^{\prime})=\phi(\pi)-\phi(\pi^{\prime})-\langle\nabla\phi(\pi^{\prime}),\pi-\pi^{\prime}\rangle\,,$
where we suppose throughout this paper that the function $\phi$ is of Legendre type, i.e. strictly convex and essentially smooth in the relative interior of $\operatorname{dom}\phi$ (see Rockafellar (1970), section 26). The mirror map choice gives rise to a large class of algorithms. Two special cases are noteworthy: (a) When $\phi$ is the squared $2$-norm, the Bregman divergence is squared Euclidean distance and the resulting algorithm is a projected $Q$-ascent and (b) when $\phi$ is the negative entropy, the corresponding Bregman divergence is the Kullback-Leibler (KL) divergence and (PMD) becomes the popular Natural Policy Gradient algorithm (see e.g. Xiao (2022), section 4 for details).
@@
where $D_{\phi}(\pi,\pi_{k})\in\mathbb{R}^{S}$ is a vector whose components are given by $D_{\phi}(\pi_{s},\pi^{k}_{s})$ for $s\in\mathcal{S}\,.$ Note that the minimization can be carried out independently for each state component.
@@ 3.2 Policy Iteration with Lookahead
As discussed in Efroni et al. (2018a), Policy Iteration can be generalized by performing $h\geq 1$ greedy policy improvement steps at each iteration instead of a single step. The resulting $h$-PI update rule can be written:
@@
where $h\in\mathbb{N}\setminus\{0\}\,.$ The $h$-greedy policy $\pi^{k+1}$ w.r.t $V^{\pi_{k}}$ selects the first optimal action of a non-stationary $h$-horizon optimal control problem. It can also be seen as the $1$-step greedy policy w.r.t. ${\mathcal{T}}^{h-1}V^{\pi_{k}}\,.$ In the rest of this paper, we will denote by $\mathcal{G}_{h}(V)$ the set of $h$-step greedy policies w.r.t. any value function $V\in{{\mathbb{R}}}^{S}$, i.e. $\mathcal{G}_{h}(V):=\operatorname{argmax}_{\pi\in\Pi}{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V=\{\pi\in\Pi:{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V={\mathcal{T}}^{h}V\}$. Notice that $\mathcal{G}_{1}(V)=\mathcal{G}(V)$. Overall, the $h$-PI method alternates between finding a $h$-greedy policy and estimating the value of this policy. Interestingly, similarly to PI, a $h$-greedy policy guarantees monotonic improvement, i.e. $V^{\pi^{\prime}}\geq V^{\pi}$ component-wise for any $\pi\in\Pi,\pi^{\prime}\in\mathcal{G}_{h}(V^{\pi})\,.$ Moreover, since ${\mathcal{T}}^{h}$ is a $\gamma^{h}$-contraction, it can be shown that the sequence $(\|V^{\star}-V^{\pi_{k}}\|_{\infty})_{k}$ is contracting with coefficient $\gamma^{h}\,.$ We refer the reader to section 3 in Efroni et al. (2018a) for a more detailed discussion regarding $h$-PI.
@@
While the iteration complexity always improves with larger depth $h$ (see Theorem 3 in Efroni et al. (2018a)), each iteration becomes more computationally demanding. As mentioned in Efroni et al. (2019), when the model is known, the $h$-greedy policy can be computed with Dynamic Programming (DP) in linear time in the depth $h$. When the model is unknown, one can use model-free algorithms such as Q-learning. Another possibility is to implement a tree-search of depth $h$ starting from the root state $s$ to compute the $h$-greedy policy in deterministic MDPs. We refer to Efroni et al. (2019) for further details. Sampling-based tree search methods such as Monte Carlo Tree Search (MCTS) (see Browne et al. (2012)) can be used to circumvent the exponential complexity in the depth $h$ of tree search methods.
@@ 4 Policy Mirror Descent with Lookahead
In this section we propose our novel Policy Mirror Descent (PMD) algorithm which incorporates $h$-step lookahead for policy improvement: $h$-PMD. Interestingly, this algorithm recovers the standard Policy Mirror Descent (PMD) (Xiao, 2022; Li et al., 2023) when $h=1$. Before introducing our algorithm, we start by describing the standard PMD algorithm and its connection with PI. Then we briefly discuss how multi-step lookahead has been incorporated to PI to motivate and present our novel $h$-PMD algorithm.
@@ 4.1 $h$-PMD: Using $h$-Step Greedy Updates
Notice that we recover (1), i.e. (PMD), by setting $h=1\,.$ We now provide an alternative way to write the $h$-PMD update rule which is more convenient for its implementation. We introduce a few additional notations for this purpose. For any policy $\pi\in\Pi,$ let $V_{h}^{\pi}:={\mathcal{T}}^{h-1}V^{\pi}$ for any integer $h\geq~{}1\,.$ Consider the corresponding state-action value function defined for every $s,a\in\mathcal{S}\times\mathcal{A}$ by:
@@
for every $s\in\mathcal{S}$ and $\text{Proj}_{\Delta(\mathcal{A})}$ is the projection operator on the simplex $\Delta(A)\,.$
@@
$\pi_{k+1}(a|s)=\pi_{k}(a|s)\frac{\exp(\eta_{k}Q_{h}^{\pi_{k}}(s,a)}{Z_{k}(s)}\,,$ (4)
where $Z_{k}(s):=\sum_{a\in\mathcal{A}}\pi_{k}(a|s)\exp(\eta_{k}Q_{h}^{\pi_{k}}(s,a))\,.$
@@ 4.2 Convergence Analysis for Exact $h$-PMD
Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in Johnson et al. (2023). In this section, we establish a linear convergence rate for the subobtimality value function gap of the exact $h$-PMD algorithm with a contraction factor of $\gamma^{h}$ where $h$ is the lookahead depth. We assume for now that the value function $V^{\pi_{k}}$ and a greedy policy $\pi_{k+1}$ in ($h$-PMD) can be computed exactly. These assumptions will be relaxed in a subsequent theorem dealing with inexact $h$-PMD.
@@
  • Similarly to Johnson et al. (2023) and unlike prior work (Xiao, 2022; Lan, 2023) featuring distribution mismatch coefficients which can scale with the size of the state space, our convergence rate does not depend on any other instance-dependent quantities.
  • In Efroni et al. (2018a), it is shown that $h$-PI also achieves this same convergence rate, but without the additional term containing the errors due to $c_{k}$. This result can be recovered in $h$-PMD by setting the stepsize $\eta_{k}$ to infinity, effectively ignoring the effect of the divergence $D_{\phi}$. However, we have found experimentally that in the inexact case this leads to more instability, and a higher suboptimality gap at convergence.
  • As noted in Efroni et al. (2018a), this faster rate comes at the cost of a more complex value function computation at each iteration. However, our experiments show that the benefits of the faster convergence rate greatly outweigh the extra cost of computing the lookahead, in terms of both overall running time until convergence and sample complexity (in the inexact case).
  • The cost of computing the adaptive step sizes might seem high due to the minimisation over greedy policies. However as noted in Johnson et al. (2023) there is typically only one greedy policy at each iteration, and in the case that there is more than one it is possible to just pick any arbitrary greedy policy and the lower bound will still be satisfied.
  • We defer the proof of Theorem 4.1 to Appendix A. Our proof highlights the relationship between PMD and PI through the explicit use of Bellman operators. Notice that even in the particular case of $h=1$, our proof is more compact compared to the one in Johnson et al. (2023) (see sec. 6 and Lemma A.1 to A.3 therein for comparison) using our convenient notations.
5 Inexact $h$-Policy Mirror Descent

### v2

@@ 1 Introduction
Policy Mirror Descent (PMD) is a general class of algorithms for solving reinforcement learning (RL) problems. Motivated by the surge of interest in understanding popular Policy Gradient (PG) methods, PMD has been recently investigated in a line of works (Lan, 2023; Xiao, 2022; Li et al., 2023; Johnson et al., 2023). Notably, PMD encompasses several PG methods as particular cases via its flexible mirror map. A prominent example is the celebrated Natural PG method. PMD has also close connections to state-of-the-art methods such as TRPO and PPO (Tomar et al., 2022) which have achieved widespread empirical success (Schulman et al., 2015, 2017), including most recently in fine-tuning Large Language Models via RL from human feedback (Ouyang et al., 2022). Interestingly, PMD has also been inspired by one of the most fundamental algorithms in RL: Policy Iteration (PI). While PI alternates between policy evaluation and policy improvement, PMD regularizes the latter step to address the instability issue of PI with inexact policy evaluation.
@@
  • We examine the inexact setting where the $h$-step action value functions are not available. In this setting, we propose a Monte Carlo sample-based procedure to estimate the unknown $h$-step lookahead action-value function at each state-action pair. We provide a sample complexity for this Monte Carlo procedure, improving over standard PMD thanks to the use of lookahead. Larger lookahead depth translates into a better sample complexity and a faster suboptimality gap convergence rate. However, this improvement comes at a more intensive computational effort to perform the lookahead steps. Sec. 5 discusses this inexact setting and the aforementioned tradeoff.
  • We extend our results to the function approximation setting to address the case of large state spaces where tabular methods are not tractable. In particular, we design a $h$-PMD algorithm where the $h$-step lookahead action-value function is approximated using a linear combination of state-action feature vectors in the policy evaluation step. Under linear function approximation and using a generative sampling model, we provide a performance bound for our $h$-PMD algorithm with linear function approximation. Our resulting sample complexity only depends on the dimension of the feature map space instead of the size of the state space and improves over prior work analyzing PMD with function approximation.
@@ 2 Preliminaries
Notation. For any integer $n\in\mathbb{N}\setminus\{0\}$, the set of integers from $1$ to $n$ is denoted by $[n]$. For a finite set $\mathcal{X},$ we denote by $\Delta(\mathcal{X})$ the probability simplex over the set $\mathcal{X}\,.$ For any $d\in\mathbb{N}$, we endow the Euclidean space ${{\mathbb{R}}}^{d}$ with the norm $\|\cdot\|_{\infty}$ defined for every $v\in{{\mathbb{R}}}^{d}$ by $\|v\|_{\infty}:=\max_{i\in[d]}|v_{i}|$ where $v_{i}$ is the $i$-th coordinate. The relative interior of a set $\mathcal{Z}$ is denoted by $\operatorname{ri}(\mathcal{Z})\,.$
@@
The Bellman optimality operator ${\mathcal{T}}$ and the expected Bellman operator ${\mathcal{T}^{\pi}}$ for any given policy $\pi\in\Pi$ are $\gamma$-contraction mappings w.r.t. the $\|\cdot\|_{\infty}$ norm and hence admit unique fixed points $V^{\star}$ and $V^{\pi}$ respectively. In particular, these operators satisfy the following inequalities: $\lVert{\mathcal{T}^{\pi}}V-V^{\pi}\rVert_{\infty}\leq\gamma\lVert V-V^{\pi}\rVert_{\infty}$ and $\lVert{\mathcal{T}}V-V^{\star}\rVert_{\infty}\leq\gamma\lVert V-V^{\star}\rVert_{\infty}$ for any $V\in{{\mathbb{R}}}^{S}\,.$ Moreover, we recall that for $\mathbf{e}=(1,1,\ldots,1)\in\mathbb{R}^{S}$ and any $c\in\mathbb{R}_{\geq 0},V\in{{\mathbb{R}}}^{S}$, we have ${\mathcal{T}^{\pi}}(V+c\,\mathbf{e})={\mathcal{T}^{\pi}}V+\gamma c\,\mathbf{e}$ and ${\mathcal{T}}(V+c\,\mathbf{e})={\mathcal{T}}V+\gamma c\,\mathbf{e}\,.$ We recall that the set $\mathcal{G}(V^{\star}):=\{\pi\in\Pi:{\mathcal{T}^{\pi}}V^{\star}={\mathcal{T}}V^{\star}\}$ of $1$step greedy policies w.r.t. the optimal value $V^{\star}$ coincides with the set of stationary optimal policies.
@@ 3.1 PMD and its Connection to PI
Policy Gradient methods consist in performing gradient ascent updates with respect to an objective $\mathbb{E}_{\rho}[V^{\pi}(s)]$ w.r.t. the policy $\pi$ parametrized by its values $\pi(a|s)$ for $s,a\in\mathcal{S}\times\mathcal{A}$ in the case of a tabular policy parametrization. Interestingly, policy gradient methods have been recently shown to be closely connected to a class of Policy Mirror Descent algorithms by using dynamical reweighting of the Bregman divergence with discounted visitation distribution coefficients in a classical mirror descent update rule using policy gradients. We refer the reader to section 4 in Xiao (2022) or section 3.1 in Johnson et al. (2023) for a detailed exposition. The resulting PMD update rule is given by
@@
where we use the shorthand notations $\pi^{k}_{s}={\pi_{k}}(\cdot|s)$ and $\pi_{s}=\pi(\cdot|s)$, where $(\eta_{k})$ is a sequence of positive stepsizes, $Q^{\pi_{k}}_{s}\in\mathbb{R}^{A}$ is a vector containing state action values for ${\pi_{k}}$ at state $s$, and $D_{\phi}$ is the Bregman divergence induced by a mirror map $\phi:\operatorname{dom}\phi\to{{\mathbb{R}}}$ such that $\Delta(\mathcal{A})\subset\operatorname{dom}\phi\,,$ i.e. for any $\pi,\pi^{\prime}\in\Pi,s\in\mathcal{S},$
$D_{\phi}(\pi_{s},\pi^{\prime}_{s})=\phi(\pi_{s})-\phi(\pi^{\prime}_{s})-\langle\nabla\phi(\pi^{\prime}_{s}),\pi_{s}-\pi^{\prime}_{s}\rangle\,,$
where we suppose throughout this paper that the function $\phi$ is of Legendre type, i.e. strictly convex and essentially smooth in the relative interior of $\operatorname{dom}\phi$ (see Rockafellar (1970), section 26). The mirror map choice gives rise to a large class of algorithms. Two special cases are noteworthy: (a) When $\phi$ is the squared $2$-norm, the Bregman divergence is the squared Euclidean distance and the resulting algorithm is a projected $Q$-ascent and (b) when $\phi$ is the negative entropy, the corresponding Bregman divergence is the Kullback-Leibler (KL) divergence and (PMD) becomes the popular Natural Policy Gradient algorithm (see e.g. Xiao (2022), section 4 for details).
@@
where $D_{\phi}(\pi,\pi_{k})\in\mathbb{R}^{S}$ is a vector whose components are given by $D_{\phi}(\pi_{s},\pi^{k}_{s})$ for $s\in\mathcal{S}\,.$ Note that the maximization can be carried out independently for each state component.
@@ 3.2 Policy Iteration with Lookahead
As discussed in Efroni et al. (2018a), Policy Iteration can be generalized by performing $h\geq 1$ greedy policy improvement steps at each iteration instead of a single step. The resulting $h$-PI update rule is:
@@
where $h\in\mathbb{N}\setminus\{0\}\,.$ The $h$-greedy policy $\pi^{k+1}$ w.r.t $V^{\pi_{k}}$ selects the first optimal action of a non-stationary $h$-horizon optimal control problem. It can also be seen as the $1$-step greedy policy w.r.t. ${\mathcal{T}}^{h-1}V^{\pi_{k}}\,.$ In the rest of this paper, we will denote by $\mathcal{G}_{h}(V)$ the set of $h$-step greedy policies w.r.t. any value function $V\in{{\mathbb{R}}}^{S}$, i.e. $\mathcal{G}_{h}(V):=\operatorname{argmax}_{\pi\in\Pi}{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V=\{\pi\in\Pi:{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V={\mathcal{T}}^{h}V\}$. Notice that $\mathcal{G}_{1}(V)=\mathcal{G}(V)$. Overall, the $h$-PI method alternates between finding a $h$-greedy policy and estimating the value of this policy. Interestingly, similarly to PI, a $h$-greedy policy guarantees monotonic improvement, i.e. $V^{\pi^{\prime}}\geq V^{\pi}$ component-wise for any $\pi\in\Pi,\pi^{\prime}\in\mathcal{G}_{h}(V^{\pi})\,.$ Moreover, since ${\mathcal{T}}^{h}$ is a $\gamma^{h}$-contraction, it can be shown that the sequence $(\|V^{\star}-V^{\pi_{k}}\|_{\infty})_{k}$ is contracting with coefficient $\gamma^{h}\,.$ See section 3 in Efroni et al. (2018a) for further details about $h$-PI.
@@
While the iteration complexity always improves with larger depth $h$ (see Theorem 3 in Efroni et al. (2018a)), each iteration becomes more computationally demanding. As mentioned in Efroni et al. (2019), when the model is known, the $h$-greedy policy can be computed with Dynamic Programming (DP) in linear time in the depth $h$.
Another possibility is to implement a tree-search of depth $h$ starting from the root state $s$ to compute the $h$-greedy policy in deterministic MDPs. We refer to Efroni et al. (2019) for further details. Sampling-based tree search methods such as Monte Carlo Tree Search (MCTS) (see Browne et al. (2012)) can be used to circumvent the exponential complexity in the depth $h$ of tree search methods.
@@ 4 Policy Mirror Descent with Lookahead
In this section we propose our novel Policy Mirror Descent (PMD) algorithm which incorporates $h$-step lookahead for policy improvement: $h$-PMD.
@@ 4.1 $h$-PMD: Using $h$-Step Greedy Updates
Notice that we recover (1), i.e. (PMD), by setting $h=1\,.$ We now provide an alternative way to write the $h$-PMD update rule which is more convenient for its implementation. We introduce a few additional notations for this purpose. For any policy $\pi\in\Pi,$ let $V_{h}^{\pi}:={\mathcal{T}}^{h-1}V^{\pi}$ for any integer $h\geq~{}1\,.$ Consider the corresponding action value function defined for every $s,a\in\mathcal{S}\times\mathcal{A}$ by:
@@
for every $s\in\mathcal{S}$ and $\text{Proj}_{\Delta(\mathcal{A})}$ is the projection operator on the simplex $\Delta(\mathcal{A})\,.$
@@
$\pi_{k+1}(a|s)=\pi_{k}(a|s)\frac{\exp(\eta_{k}Q_{h}^{\pi_{k}}(s,a))}{Z_{k}(s)}\,,\quad Z_{k}(s):=\sum_{a\in\mathcal{A}}\pi_{k}(a|s)\exp(\eta_{k}Q_{h}^{\pi_{k}}(s,a))\,.$ (4)
Connection to AlphaZero. Before switching gears to the analysis, we point out an interesting connection between $h$-PMD and the successful class of AlphaZero algorithms Silver et al. (2017, 2018); Schrittwieser et al. (2019) which are based on heuristics.
It has been shown in Grill et al. (2020b) that AlphaZero can be seen as a regularized policy optimization algorithm, drawing a connection to the standard PMD algorithm (with $h=1$). We argue that AlphaZero is even more naturally connected to our $h$-PMD algorithm. Indeed, the $Q$ values in (Grill et al., 2020b, Eq. (7) p. 3) correspond more closely to the lookahead action value function $Q_{h}^{\pi}$ approximated via tree search instead of the standard $Q^{\pi}$ function with $h=1$. This connection clearly delineates the dependence on the lookahead depth (or tree search depth) $h$ which was not clear in Grill et al. (2020b). We have implemented a version of our $h$-PMD algorithm with Deep Mind’s MCTS implementation (see section D.6 and the code provided for details). Conducting further experiments to compare our algorithm to AlphaZero on similar large scale settings would be interesting. We are not aware of any theoretical convergence guarantee for AlphaZero which relies on many heuristics. In contrast, our $h$-PMD algorithm enjoys theoretical guarantees that we establish in the next sections.
@@ 4.2 Convergence Analysis for Exact $h$-PMD
Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in Johnson et al. (2023). In this section, we establish a linear convergence rate for the subobtimality value function gap of the exact $h$-PMD algorithm with a contraction factor of $\gamma^{h}$ where $h$ is the lookahead depth. We assume for now that the value function $V^{\pi_{k}}$ and a greedy policy $\pi_{k+1}$ in ($h$-PMD) can be computed exactly. These assumptions will be relaxed in the next section dealing with inexact $h$-PMD.
@@
  • Unlike prior work (Xiao, 2022; Lan, 2023) featuring distribution mismatch coefficients which can scale with the size of the state space, our convergence rate does not depend on any other instance-dependent quantities. This is thanks to the analysis which circumvents the use of the performance difference lemma similarly to Johnson et al. (2023).
  • Choosing $c_{t}=\gamma^{2h(t+1)}c_{0}$ for a positive constant $c_{0}$ for every integer $t$, the bound becomes (after bounding the geometric sum): $\lVert V^{\star}-V^{\pi_{k}}\rVert_{\infty}\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{c_{0}}{(1-\gamma)^{2}}\right)$. As $c_{0}\rightarrow 0$ we recover the linear convergence result of $h$-PI established in Efroni et al. (2018a).
  • This faster rate comes at the cost of a more complex value function computation at each iteration. However, our experiments suggest that the benefits of the faster convergence rate greatly outweigh the extra cost of computing the lookahead, in terms of both overall running time until convergence and sample complexity (in the inexact case). See section 7 and Appendix D for evidence.
  • The cost of computing the adaptive stepsizes is typically simply that of computing a Bregman divergence between two policies. See Appendix A.3, D.4 for a more detailed discussion.
  • We defer the proof of Theorem 4.1 to Appendix A. Our proof highlights the relationship between $h$-PMD and $h$-PI through the explicit use of Bellman operators. Notice that even in the particular case of $h=1$, our proof is more compact compared to the one in Johnson et al. (2023) (see sec. 6 and Lemma A.1 to A.3 therein for comparison) using our convenient notations.
5 Inexact and Stochastic $h$-Policy Mirror Descent